# A Survey of Artificial Intelligence in Gait-Based Neurodegenerative Disease Diagnosis

### v2

@@ Abstract
Recent years have witnessed an increasing global population affected by neurodegenerative diseases (NDs), which traditionally require extensive healthcare resources and human effort for medical diagnosis and monitoring. As a crucial disease-related motor symptom, human gait can be exploited to characterize different NDs. The current advances in artificial intelligence (AI) models enable automatic gait analysis for NDs identification and classification, opening a new avenue to facilitate faster and more cost-effective diagnosis of NDs. In this paper, we provide a comprehensive survey on recent progress of machine learning and deep learning based AI techniques applied to diagnosis of five typical NDs through gait. We provide an overview of the process of AI-assisted NDs diagnosis, and present a systematic taxonomy of existing gait data and AI models. Meanwhile, a novel quality evaluation criterion is proposed to quantitatively assess the quality of existing studies. Through an extensive review and analysis of 164 studies, we identify and discuss the challenges, potential solutions, and future directions in this field. Finally, we envision the prospective utilization of 3D skeleton data for human gait representation and the development of more efficient AI models for NDs diagnosis [^1] We provide a public resource repository to track and facilitate developments in this emerging field: https://github.com/Kali-Hac/AI4NDD-Survey..
@@ 1 Introduction
Neurodegenerative Diseases (NDs) such as Alzheimer’s disease (AD) and Parkinson’s disease (PD) are among the most widespread and devastating disorders affecting millions of people worldwide [1]. According to the Alzheimer’s Disease Association [2] and the Parkinson’s Foundation Report [3] in 2023, nearly 6.2 million people and 1 million people are diagnosed with AD and PD respectively in the United States. While in China, there are over 9.8 million AD patients [4], and it is estimated that Chinese PD patients will increase to 4.9 million by 2030, accounting for a half of the worldwide PD population [5]. From 1990 to 2019, the effected global population of AD, PD, and other neurodegenerative diseases ($e.g.$, amyotrophic lateral sclerosis) has significantly increased (illustrated in Fig. 2), leading to a growing burden on global healthcare system. On the one hand, these diseases could result in pathological gaits and chronic pain in body joints, tissues, and nerves [6, 7], which severely reduce the flexibility, stability, mobility, and other functional capabilities of a human body. On the other hand, they also impose large psychological and financial burdens to patients and their family [8, 9, 10]. The treatment cost for these diseases has reached approximately USD 130 billion per year and further rise has been estimated [11, 12]. Therefore, it is essential to conduct an earlier and more reliable diagnosis of these diseases, so as to provide a timely proper treatment to mitigate the burden of societal and healthcare resources.
@@
In recent years, driven by the widespread availability of body positioning technologies and economical portable devices to monitor human gait, gait-based NDs diagnosis ($i.e.,$ NDs classification, identification or prediction) has propelled a surge of attention across academic, industrial, and medical communities [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. To perform NDs diagnosis from diverse gait data, the Artificial Intelligence (AI) technologies, especially Machine Learning (ML) and Deep Learning (DL) models, have been extensively explored and applied [46, 34, 35, 47, 36, 37]. A common practice is to extract gait parameters ($e.g.,$ stride length, stance duration, gait cycle time) from sensor-based data such as vertical ground reaction force (vGRF) or vision-based data such as gait images, and leverage ML algorithms ($e.g.,$ support vector machine (SVM)) or deep neural networks to automatically recognize gait patterns associated with NDs to aid in their diagnosis [48, 49, 50]. Compared with traditional diagnosis that requires manual observations of clinical motor symptoms ($e.g.,$ rigidity), these AI-based methods can analyze the abnormality in gait to automatically identify NDs related patterns, which shows immense potential in assisting healthcare professionals to diagnose and treat these diseases more efficiently [36, 37].
In this paper, we systematically review recent advancements of utilizing AI to help diagnose NDs from human gait. In particular, our work focuses on the diagnosis of five most prevalent and typical NDs (PD, AD, ALS, Huntington’s Disease (HD), Multiple Sclerosis (MS)) and all existing AI models that can help diagnose these diseases ($i.e.,$ identify, classify or predict NDs) from human gait data. A total of 2069 potentially relevant articles, including 14 survey papers, are found from three databases (PubMed, Web of Science (WoS), Google Scholar). We first screen these articles by titles and abstracts based on the pre-defined inclusion criteria, and then conduct a full-text assessment to exclude irrelevant papers. Finally, 164 of 2069 papers are selected to be included in our survey.
Figure 3: Different phases of normal human gait with coordinated movements of legs and arms. Right: Examples for hemiplegic gait.
Figure 4: Examples for hemiplegic gait (left), Parkinsonian gait (middle), and neuropathic gait (right) [51].
@@ 1.1 Objectives and Contributions
The objectives of this survey can be encapsulated in the following six key questions (Q): (Q1) How is human gait associated with NDs? (Q2) What is the process of AI-assisted NDs diagnosis from gait data? (Q3) What are various gait data types, gait collecting approaches, and AI model types for NDs diagnosis? (Q4) What are state-of-the-art AI models and their performance results ($e.g.,$ diagnostic accuracy) in gait-based NDs diagnosis? (Q5) What are the current main challenges and promising future directions of this field? (Q6) Is it feasible to exploit 3D skeleton data to represent human gait and build AI models for NDs diagnosis?
The contributions of this work are as follows. We provide a systematic literature review of existing AI technologies for gait-based diagnosis of five most widespread NDs (PD, AD, ALS, HD, MS). We formulate the general process of AI-assisted gait-based NDs diagnosis, and present a systematic taxonomy of used gait data and AI models. Our survey demonstrates how human gait features can be a crucial indicator for diagnosing various NDs, and highlights the significant role of current AI models in facilitating the automation of this diagnostic process. Moreover, we analyze the statistics of existing studies to reveal the current development and future trend of this area, and comprehensively discuss the key challenges in utilizing AI for gait-based NDs diagnosis, such as scarcity and imbalance of gait data, lack of multiple data modalities and sources, and unreliable AI model designs ($e.g.,$ limited generalizability, efficiency, and interpretability). We elaborate on these issues and delve further into their potential solutions, related technologies, and future research directions. To the best of our knowledge, this is the most elaborated survey on this topic, and it also provides a practical roadmap for future investigations of AI technologies for NDs diagnosis. As our research vision, 3D skeleton data, which is an emerging generic and scalable data modality, is for the first time explored to characterize gait patterns associated with NDs. We empirically present a 3D skeleton based framework for NDs diagnosis, paving the way for more efficient gait representations and AI model learning in this field.
@@ 1.2 Organization
Aligning with the above objectives, we organize the rest of this survey as follows. Sec. 2 provides the definitions of fundamental concepts related to gait, NDs, and gait-based NDs diagnosis (Q1). It also presents an overview for AI-assisted gait-based NDs diagnosis process (Q2), and a systematic taxonomy of gait data types, gait collection technologies, and AI model types (Q3). Sec. 3 elaborates on the literature screening result with a multi-faceted statistics analysis, and provide a content summary of included studies (Q4). In Sec. 5, we discuss the main challenges, potential solutions, and several promising directions in this area (Q5), while proposing our research vision on exploiting 3D skeleton data for NDs diagnosis (Q6). Finally, we conclude this survey in Sec. 6.
@@ 2.1 Basic Concepts
    Neurodegenerative Diseases (NDs): The diseases with progressive loss of nerve cells, neuron structure, or/and their functions in the brain or peripheral nervous system are collectively termed NDs. The focus of this study is on the five most common NDs, including AD, PD, ALS, HD, and MS.
@@ 2.2.1 Neurodegenerative Disease Related Gaits
  • Neuropathic gait is characterized by foot drop, where one foot flops down when lifting the leg, necessitating a higher lifting of the knee to prevent the toe from dragging on the ground during walking (see Fig. 4). Neuropathic gait may be a symptom of amyotrophic lateral sclerosis (ALS), multiple sclerosis (MS) or peripheral neuropathy.
@@
  • Choreiform gait is characterized by the irregular, jerky, involuntary movements in all extremities. This gait can be seen in certain basal ganglia disorders including Huntington’s Disease (HD), Sydenham’s chorea, and other forms of chorea, athetosis or dystonia.
@@ 2.4 Gait Data Taxonomy
  • Sensor Modality: Advancement in sensor technologies allows its application in various tasks such as gait analysis, action recognition, etc. The sensors explored among the collected papers offer various advantages and could be classified into these four types, including force, inertial, hybrid, and other less prevalent sensors [57, 58, 59, 60, 41].
@@
      Force Sensor (FS): There are mainly two types of force or pressure capturing approaches. The first is to use non-wearable floor-mounted devices ($e.g$, force mats, treadmill) to measure vertical Ground Reaction Force (vGRF), moments, or plantar pressure profiles [41]. In a laboratory setting, these stationary floor-mounted devices are sufficient for precise data collection. However, since those devices are often bulky, operating-costly, and expensive, it’s not often clinically applicable. Therefore, researchers develop some small wearable devices ($e.g.$, insole) whose application in the field has became a popular choice and risen significantly in the past decade, due to its portable features and cheap installment, at the cost of less accuracy and reliability compare with the floor-mounted ones [61, 62].
@@
      Inertial Sensor: Also termed as Inertial Measurement Unit (IMU), is often attached to limb segments ($e.g$, foot) [63] or embedded in wearable devices ($e.g.$, smartwatch) to capture dynamic motion ($e.g.$, 3D-acceleration) of the subject [64, 65]. These sensors or devices record the repeating pattern of gait signals generated by the locomotion of the subjects. These signals will then be interpreted with various algorithms for the intended tasks, including gait event detection, kinetic or kinematic parameters estimation, and gait classification.
@@
      Multi-Sensor: These sensors typically combine various types of sensors for a more comprehensive data collection. Among the surveyed papers, studies such as [66, 67, 68] utilize both IMU sensors and Force Sensitive Resistors (FSRs), while [69] employs Electroencephalogram (EEG) sensors and FSRs. These combinations harvest the strengths of each sensor type hence help enhancing the reliability and accuracy of gait event detection and classification.
@@
      Other Sensors: Sensors included in this category are less prevalent and exploited among collected papers. However, these sensors still play an important role in providing a more full-sided understanding of the human gait. For instance, EEG sensor is used to measure the brain electrical activity signals data which can be used to analyze brain activity patterns that may correlate with specific gait abnormalities [70, 69, 71]. Electromyography (EMG) sensor is another useful sensor which is used to measure the electrical activity produced by muscles whose activities are controlled by the motor neurons [72]. Since NDs patients suffer from the degeneration of nerve cells, they manifest significantly different muscle electrical pattern from healthy subjects [36]. Therefore, their EMG data can be utilized to aid NDs diagnosis [72].
  • Vision Modality: The vision modality employs optoelectronic Motion capture (Mocap)systems, including various types of cameras ($e.g.$, Infrared (IR) Camera [73]), to capture the locomotion of subjects during walking. It’s used to overcome the limitations of the traditional clinical examination which purely depends on the physicians’ naked eyes. Depends on the precision requirement of the video data, the vision modality could be further divided into two sub-categories, namely marker-based and marker-less systems [47].
@@
      Marker-based: This approach is developed at an early stage when the cameras are not advanced enough for precise motion capture. Therefore, these Mocap systems ($e.g.$, Vicon) use reflective markers to help capture targets’ movement in the images more precisely. The advantages of using the marker-based approach are that they’re highly accurate and provides exact location of the body landmark. These advantages hence make this modality a gold standard in the field. However, this method requires attaching markers to the body, leading to additional expenses and the potential of alter the gait [74, 75, 76].
@@
      Marker-free: Due to the limitations of the marker-based approach and the development of the camera technology, the marker-free approach comes into appearance. This approach derives corresponding features from the target’s gait video without using any markers. Another highlight is the emerging application of smartphones in this field, which, despite being in its infancy, offers new insights due to its advantages like compactness, portability, and affordability. However, it still faces challenges such as lower accuracy in tracking movements of body parts compared to the gold standard ($i.e.,$ marker-based approaches).
  • Sensor+Vision Modality: This modality could also be referred to as multi-modal modality. In comparison with single-modal approaches, although the latter ones have their own merits and show some promising results, this combined approach unleashes a bigger potential, such as providing richer data and more comprehensive features, capturing more subtle changes in gait, improving recognition performance [71, 77, 78]. In particular, the application of Microsoft Kinect system has gain a tremendous amount of attention in the field due to its utilization of both depth sensor and RGB camera. This integration enables a more accurate estimation of the gait parameters without sacrificing the advantages mentioned in the marker-less vision modality. More specifically, using conventional cameras to estimate skeleton coordinates produces coarse results. However, when employing the depth sensor, the spatiotemporal features can be precisely derived from the skeleton joints’ coordinates obtained using it. The precision and efficacy of using Kinect has been verified in many studies which compare the results with the gold standard (marker-based vision) approach. Nevertheless, one limitation of current Kinect system is that its precision reduces as the subject walks further away from the camera, resulting in a width and length-limited data collection space [79, 80, 81, 82]. Besides the RGB-D camera approach, there are also other combinations of these two modalities. For example, Tahir et al. fused a marker-based IR camera and a force-sensitive platform [83], while Chatzaki et al. utilized a conventional camera, IMU, and a force platform [78]. Additionally, Zhao et al. merge some vGRF and IMU sensors with Kinect [77].

### v3

@@ Abstract
Recent years have witnessed an increasing global population affected by neurodegenerative diseases (NDs), which traditionally require extensive healthcare resources and human effort for medical diagnosis and monitoring. As a crucial disease-related motor symptom, human gait can be exploited to characterize different NDs. The current advances in artificial intelligence (AI) models enable automatic gait analysis for NDs identification and classification, opening a new avenue to facilitate faster and more cost-effective diagnosis of NDs. In this paper, we provide a comprehensive survey on recent progress of machine learning and deep learning based AI techniques applied to diagnosis of five typical NDs through gait. We provide an overview of the process of AI-assisted NDs diagnosis, and present a systematic taxonomy of existing gait data and AI models. Meanwhile, a novel quality evaluation criterion is proposed to quantitatively assess the quality of existing studies. Through an extensive review and analysis of 169 studies, we present recent technical advancements, discuss existing challenges, potential solutions, and future directions in this field. Finally, we envision the prospective utilization of 3D skeleton data for human gait representation and the development of more efficient AI models for NDs diagnosis [^1] We provide a public resource repository to track and facilitate developments in this emerging field: https://github.com/minlinzeng/AI4NDD-Survey..
@@ 1 Introduction
Neurodegenerative Diseases (NDs) such as Alzheimer’s disease (AD) and Parkinson’s disease (PD) are among the most widespread and devastating disorders affecting millions of people worldwide [1]. According to the Alzheimer’s Disease Association [2] and the Parkinson’s Foundation Report [3] in 2023, nearly 6.2 million people and 1 million people are diagnosed with AD and PD in the United States, respectively. While in China, there are over 9.8 million AD patients [4], and it is estimated that Chinese PD patients will increase to 4.9 million by 2030, accounting for a half of the worldwide PD population [5]. From 1990 to 2019, the effected global population of AD, PD, and other neurodegenerative diseases ($e.g.$, amyotrophic lateral sclerosis) has significantly increased (illustrated in Fig. 2), leading to a growing burden on global healthcare system. On the one hand, these diseases could result in pathological gaits and chronic pain in body joints, tissues, and nerves [6, 7], which severely reduce the flexibility, stability, mobility, and other functional capabilities of a human body. On the other hand, they also impose large psychological and financial burdens to patients and their family [8, 9, 10]. The treatment cost for these diseases has reached approximately USD 130 billion per year and further rise has been estimated [11, 12]. Therefore, it is essential to conduct an earlier and more reliable diagnosis of these diseases, so as to provide a timely proper treatment to mitigate the burden of societal and healthcare resources.
@@
In recent years, driven by the widespread availability of body positioning technologies and economical portable devices to monitor human gait, gait-based NDs diagnosis ($i.e.,$ NDs classification, identification or prediction) has propelled a surge of attention across academic, industrial, and medical communities [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. To perform NDs diagnosis from diverse gait data, the Artificial Intelligence (AI) technologies, especially Machine Learning (ML) and Deep Learning (DL) models, have been extensively explored and applied [46, 34, 35, 47, 36, 37, 48, 49, 50, 51, 52]. A common practice is to extract gait parameters ($e.g.,$ stride length, stance duration, gait cycle time) from sensor-based data such as vertical ground reaction force (vGRF) or vision-based data such as gait images, and leverage ML algorithms ($e.g.,$ support vector machine (SVM)) or deep neural networks to automatically recognize gait patterns associated with NDs to aid in their diagnosis [53, 54, 55]. Compared with traditional diagnosis that requires manual observations of clinical motor symptoms ($e.g.,$ rigidity), these AI-based methods can analyze the abnormality in gait to automatically identify NDs related patterns, which shows immense potential in assisting healthcare professionals to diagnose and treat these diseases more efficiently [36, 37].
In this paper, we systematically review recent advancements of utilizing AI to help diagnose NDs from human gait. In particular, our work focuses on the diagnosis of five most prevalent and typical NDs (PD, AD, ALS, Huntington’s Disease (HD), Multiple Sclerosis (MS)) and all existing AI models that can help diagnose these diseases ($i.e.,$ identify, classify or predict NDs) from human gait data. A total of 2074 potentially relevant articles, including 14 survey papers, are found from three databases (PubMed, Web of Science (WoS), Google Scholar). We first screen these articles by titles and abstracts based on the pre-defined inclusion criteria, and then conduct a full-text assessment to exclude irrelevant papers. Finally, 169 of 2074 papers are selected to be included in our survey.
Figure 3: Different phases of normal human gait with coordinated movements of legs and arms.
Figure 4: Examples for hemiplegic gait (left), Parkinsonian gait (middle), neuropathic gait (right) [56].
@@ 1.1 Objectives and Contributions
The objectives of this survey can be encapsulated in the following six key questions (Q):
  • Q1. How is human gait associated with NDs?
  • Q2. What is the process of AI-assisted NDs diagnosis from gait data?
  • Q3. What are various gait data types, gait collecting approaches, and AI model types for NDs diagnosis?
  • Q4. What are different AI models used in the gait-based diagnosis of different NDs, and how are their performance ($e.g.,$ diagnostic accuracy) and quality?
  • Q5. What are the most representative datasets and state-of-the-art AI models in gait-based NDs diagnosis?
  • Q6. What are the current main challenges and promising future directions of this field?
  • Q7. Is it feasible to exploit 3D skeleton data to represent human gait and build AI models for NDs diagnosis?
The contributions of this work are as follows. We provide a systematic literature review of existing AI technologies for gait-based diagnosis of five most widespread NDs (PD, AD, ALS, HD, MS). We formulate the general process of AI-assisted gait-based NDs diagnosis, and present a systematic taxonomy of used gait data and AI models. Our survey demonstrates how human gait features can be a crucial indicator for diagnosing various NDs, and highlights the significant role of current AI models in facilitating the automation of this diagnostic process. Moreover, we analyze the statistics of existing studies to reveal the current development and future trend of this area, and comprehensively discuss the key challenges in utilizing AI for gait-based NDs diagnosis, such as scarcity and imbalance of gait data, lack of multiple data modalities and sources, and unreliable AI model designs ($e.g.,$ limited generalizability, efficiency, and interpretability). We elaborate on these issues and delve further into their potential solutions, related technologies, and future research directions. In addition, our work also introduces a systematic quality evaluation criterion, focusing on AI novelty, comparison comprehensiveness, and sample sufficiency, to facilitate the assessment and comparison of different AI-based studies. To the best of our knowledge, this is the most elaborated survey on this topic, and it also provides a practical roadmap for future investigations of AI technologies for NDs diagnosis. As our research vision, 3D skeleton data, which is an emerging generic and scalable data modality, is for the first time explored to characterize gait patterns associated with NDs. We empirically present a 3D skeleton based framework for NDs diagnosis, paving the way for more efficient gait representations and AI model learning in this field.
@@ 1.2 Organization
Aligning with the above objectives, we organize the rest of this survey as follows. Sec. 2 provides the definitions of fundamental concepts related to gait, NDs, and gait-based NDs diagnosis (Q1). It also presents an overview for AI-assisted gait-based NDs diagnosis process (Q2), and a systematic taxonomy of gait data types, gait collection technologies, and AI model types (Q3). Sec. 3 and 4 elaborate on the literature screening result with a multi-faceted statistics analysis, and provide a content summary of included studies with a systematic quality evaluation (Q4). Sec. 5 compares the differences between our survey and existing surveys on similar topics. Sec. 6 provides an introduction for representative public datasets and state-of-the-art methods in the gait-based diagnosis of different NDs (Q5). In Sec. 7, we discuss the main challenges, potential solutions, and several promising directions in this area (Q6), while proposing our research vision on exploiting 3D skeleton data for NDs diagnosis (Q7). Finally, we conclude this survey in Sec. 8.
@@ 2.1 Basic Concepts
    Neurodegenerative Diseases: The diseases with progressive loss of nerve cells, neuron structure, or/and their functions in the brain or peripheral nervous system are collectively termed NDs. The focus of this study is on the five most common NDs, including AD, PD, ALS, HD, and MS.
@@ 2.2.1 Neurodegenerative Disease Related Gaits
  • Neuropathic gait is characterized by foot drop, where one foot flops down when lifting the leg, necessitating a higher lifting of the knee to prevent the toe from dragging on the ground during walking (see Fig. 4). Neuropathic gait may be a symptom of ALS), multiple sclerosis (MS) or peripheral neuropathy.
@@
  • Choreiform gait is characterized by the irregular, jerky, involuntary movements in all extremities. This gait can be seen in certain basal ganglia disorders including HD, Sydenham’s chorea, and other forms of chorea, athetosis or dystonia.
@@ 2.4 Gait Data Taxonomy
  • Sensor Modality: Advancement in sensor technologies allows its application in various tasks, such as gait analysis. The sensors explored among the collected papers offer various advantages and could be classified into four types, including force, inertial, hybrid, and others [62, 63, 64, 65, 41].
@@
      Force Sensor (FS): There are mainly two types of force or pressure capturing approaches. The first is to use precise non-wearable floor-mounted devices ($e.g$, force mats, treadmill) to measure vGRF, moments, or plantar pressure profiles in a laboratory setting [41]. However, since those devices are often bulky, operating-costly, and expensive, it is not often clinically applicable. Consequently, researchers have developed some small wearable devices ($e.g.$, insole) whose application in the field has risen significantly in the past decade to combat previous limitations, however at the cost of less accuracy and reliability [66, 67].
@@
      Inertial Sensor: Also termed as Inertial Measurement Unit (IMU), is often attached to limb segments ($e.g$, foot) [68] or embedded in wearable devices ($e.g.$, smartwatch) to capture dynamic motion signal ($e.g.$, 3D-acceleration) of the subject, which will then be interpreted with various algorithms for the intended tasks, such as gait event detection [69, 70].
@@
      Multi-Sensor: This category combines various types of sensor for a more comprehensive data collection. For instance, some studies [71, 72, 73] utilize both IMU sensors and Force Sensitive Resistors (FSRs), while others [74] employs Electroencephalogram (EEG) sensors and FSRs. These combinations harvest the strengths of each sensor type to enhance the reliability and accuracy of intended tasks.
@@
      Other Sensors: Sensors included in this category are less exploited among collected papers and serve as a supplementary to other sensors. For instance, EEG and electromyography (EMG) sensors can be used to measure electrical activity signals of brain and muscle, respectively, to analyze their activity patterns that may correlate with specific gait abnormalities [75, 74, 76, 36, 77]. The rationale behind this is that NDs patients suffer from the degeneration of nerve cells, hence they manifest different brain and muscle electrical pattern from healthy subjects.
  • Vision Modality: The vision modality employs optoelectronic Motion capture (Mocap) systems, encompassing various types of cameras ($e.g.$, Infrared (IR) Camera [78]) to capture the locomotion of subjects during walking. It can be used to overcome the limitations of the traditional clinical examination (see Sec. 2.1). Depending on the precision requirement for the data, this modality could be further divided into two sub-categories, namely marker-based and marker-free systems [47].
@@
      Marker-based: These approaches are developed at an early stage when the cameras are not precise enough for motion capturing. Therefore, instead of solely relying on ($e.g.$, Vicon) reflective markers to help tracking the exact location of the body landmarks during movement. Data captured by these approaches are verified in many studies with the highest precision. Therefore, these approaches are considered as the ’gold standard’ in the field. However, the body-attached markers have the potential of altering gait. Additionally, this approach incurs high costs in terms of time, manpower, and expenses due to the complex setup [79, 80, 81].
@@
      Marker-free: Due to the limitations of the marker-based approach and the development of the camera technology, the marker-free approach comes into appearance. This approach derives corresponding features from patients’ motion videos without using any markers in the data collection process. Another highlight is the emerging application of smartphones in this field, which, despite being in its infancy, offers new insights due to its advantages like compactness, portability, and affordability. However, it still faces challenges, such as lower accuracy, in tracking movements of body parts compared to the gold standard ($i.e.,$ marker-based approaches).
  • Sensor+Vision Modality: This modality is also referred to as multi-modal modality. In comparison with single-modal approaches, this combined approach unleashes a bigger potential, such as providing more comprehensive features, capturing more subtle changes in gait, and etc. [76, 82, 83].
    Microsoft Kinect, in particular, is the most applied system in this category due to its utilization of both depth sensor and RGB camera, which achieves near optimal data capturing precision while not sacrificing the advantages provided by the marker-free vision modality. More specifically, using conventional cameras to estimate skeleton coordinates can only produce coarse results. However, by combining depth sensors, the skeleton joints’ coordinates can be precisely captured. This has been verified in many studies by comparing with the gold standard approaches. Nevertheless, current Kinect system’s precision reduces as the subject walks further away from the camera, resulting in a width and length-limited data collection space [84, 85, 86, 87].
    Besides combining RGB camera and depth sensor, there are also other ways to integrate sensor and vision modalities. For example, Tahir et al. [88] fused a marker-based IR camera and a force-sensitive platform, while Chatzaki et al. [83] utilized a conventional camera, IMU, and a force platform. Additionally, Zhao et al. [82] merged some vGRF and IMU sensors with Kinect.